If the 1-year discount factor is 0.92, the interest rate must be equal to ________.
Answer #1

Interest rate =( 1/r )- 1

= (1/0.92) -1

= 1.08695 -1

= 0.8695 i.e. 8.7%

Note : r' represent to discounting factor.